Pell Manufacturing is preparing its direct labor budget for May. Projections for the month are that 33,400 units are to be produced and that direct labor time is three hours per unit. If the labor cost per hour is $12, what is the total budgeted direct labor cost for May?
Answer #1
Concepts and reason

Budgeting: Budgeting is the process of estimating the projections for future periods and communicating the information to the users. This process helps the managers to achieve the long term and short term goals.

Budgeting process: The process of formulating the future plans of a company, coordinating the planned activities, communicating the required information, and provides the standards to assess the performance of the employees and managers and motivate them is referred to as budgeting process

Expenses: The costs borne by a company to produce and sell the goods and services to the customers are known as expenses.

Fundamentals

Direct labour budget: Estimation of direct labour hours required for the budgeted units of production is known as direct labour budget

Direct labor: Direct labor includes the wages and salaries paid to workers who helps in the manufacturing of the product.

Direct labour hours: Direct labour rates are the labour costs that are incurred directly towards production of product.

Total budgeted direct labour cost: Total budgeted direct labour cost can be calculated by multiplying the total budgeted direct labour hours with the labour cost per hour. The following formula is used for calculating total budgeted direct labour cost.

Totalbudgeteddirectlaborcost}=(Totalbudgeteddirectlaborhours×Laborcostperhour)\left. \begin{array}{l}\\{\rm{Total budgeted }}\\\\{\rm{direct labor cost}}\\\end{array} \right\}{\rm{ = }}\left( \begin{array}{l}\\{\rm{Total budgeted direct labor hours}}\\\\{\rm{ \times Labor cost per hour}}\\\end{array} \right)

Calculation of total budgeted direct labour hours is given below:

Totalbudgeteddirectlaborhours}=(Budgetedunitsofproduction×Directlaborhoursperunit)=33,400units×3hoursperunit=100,200hours\begin{array}{c}\\\left. \begin{array}{l}\\{\rm{Total budgeted }}\\\\{\rm{direct labor hours}}\\\end{array} \right\}{\rm{ = }}\left( \begin{array}{l}\\{\rm{Budgeted units of production }}\\\\{\rm{ \times Direct labor hours per unit}}\\\end{array} \right)\\\\{\rm{ = 33,400 units}} \times {\rm{ 3 hours per unit}}\\\\{\rm{ = 100,200 hours}}\\\end{array}

Therefore, the total budgeted direct labour hours is 100,200 hours.

Calculation of total budgeted direct labour cost is given below:

Totalbudgeteddirectlaborcost}=(Totalbudgeteddirectlaborhours×Laborcostperhour)=100,200hours×$12=$1,202,400\begin{array}{c}\\\left. \begin{array}{l}\\{\rm{Total budgeted }}\\\\{\rm{direct labor cost}}\\\end{array} \right\}{\rm{ = }}\left( \begin{array}{l}\\{\rm{Total budgeted direct labor hours}}\\\\{\rm{ \times Labor cost per hour}}\\\end{array} \right)\\\\{\rm{ = 100,200 hours }} \times {\rm{ \$ 12}}\\\\{\rm{ = \$ 1,202,400}}\\\end{array}

Therefore, the total budgeted direct labour cost is $1,202,400.

Ans:

Total budgeted direct labour cost is $1,202,400.